A vehicle travels at 40 km/hr and skids to a stop in 16 m. How far will the same vehicle skid if it was traveling at 80 km/hr?
Answer #1

Velocity, v = 40 x 5/18 = 11.11 m/s

a = v^2/(2d)

a = 11.11^2/(2 x 16) = 3.86 m/s^2

Now again

D = ( 80* 5/18)^2/ (2 x 3.86)

D = 5.76 m

Comment in case any doubt please rate my answer....